# Regressive Side Effects of Training Language Models to Mimic Student Misconceptions

### v1

@@ Abstract
This paper presents a novel exploration into the regressive side effects of training Large Language Models (LLMs) to mimic student misconceptions for personalized education. We highlight the problem that as LLMs are trained to more accurately mimic student misconceptions, there is a compromise in the factual integrity and reasoning ability of the models. Our work involved training an LLM on a student-tutor dialogue dataset to predict student responses. The results demonstrated a decrease in the model’s performance across multiple benchmark datasets, including the ARC reasoning challenge and TruthfulQA, which evaluates the truthfulness of model’s generated responses. Furthermore, the HaluEval Dial dataset, used for hallucination detection, and MemoTrap, a memory-based task dataset, also reported a decline in the model accuracy. To combat these side effects, we introduced a “hallucination token” technique. This token, appended at the beginning of each student response during training, instructs the model to switch between mimicking student misconceptions and providing factually accurate responses. Despite the significant improvement across all datasets, the technique does not completely restore the LLM’s baseline performance, indicating the need for further research in this area. This paper contributes to the ongoing discussion on the use of LLMs for student modeling, emphasizing the need for a balance between personalized education and factual accuracy.
@@ 1 Introduction
Personalized education, an approach that caters to the unique learning needs of individuals, is increasingly becoming a key aspiration in educational technology [1, 2, 3]. With the advent of advanced Large Language Models (LLMs), this aspiration is inching closer to reality. LLMs, such as Llama [4] and GPT [5, 6] models, are playing a pivotal role in this domain, demonstrating significant potentials in various applications, including the simulation of student behavior [7] and learning patterns [8]. However, the road to leveraging LLMs for personalized education is challenging [9, 10, 11].
In this paper, we have identified regressive side effects arising from training LLMs to mimic student behavior. We find that training LLMs to replicate student misconceptions accurately is a double-edged sword. On one hand, it creates a model that can mimic student behavior, making it a potentially effective tool for personalized learning. On the other hand, it significantly compromises the model’s factual integrity and reasoning ability. These regressive side effects are a critical issue, as the primary role of any educational model is to provide accurate and reliable information.
Figure 1: This figure illustrates a typical student-tutor interaction from the CLASS [12] dataset. Unlike the CLASS methodology, which focuses on training a tutor model, our study trains a ‘student model’, with the LLM predicting student responses. This approach is motivated by the potential of personalized education, where understanding and mimicking student behavior can lead to more effective learning interventions. However, this approach, while effectively mimicking student misconceptions, leads to regressive side effects such as compromising the model’s factual integrity and diminishing its reasoning abilities. The conversation shown here exemplifies this issue, where the student’s response, while partially correct, contains misconceptions. To mitigate these side effects, we introduce hallucination tokens ([hal] and [/hal]) appended to student responses during training. These tokens instruct the model to switch between mimicking student misconceptions and providing factually accurate responses. Despite significant improvements, the technique does not fully restore the model’s baseline performance, highlighting the complexity of the issue and the need for further research.
To investigate this issue further, we conducted a comprehensive exploration involving training LLMs on a student-tutor dialogue dataset. This dataset, derived from the CLASS framework [12, 13, 2], comprises dialogues on biology questions sourced from college-level textbooks. An example of the student-tutor interaction from the dataset is illustrated in figure 1. It provides a realistic representation of student learning patterns, featuring student misconceptions and the tutor’s rectifications.
We used the dataset to train the latest Vicuna models (7B and 13B) [14], state-of-the-art Llama [4] variants, to mimic student responses. However, the training process significantly decreased the model’s performance across various benchmark datasets, including the ARC reasoning challenge [15], TruthfulQA [16], Hallucination Evaluation Dialogue [17], and MemoTrap [18]. We present a detailed analysis across nine key benchmarks using the Eleuther AI Language Model Evaluation Harness [19], a widely used framework [20] which provides a thorough and fair assessment of generative models across a spectrum of reasoning and general knowledge tasks. We present a detailed analysis across nine key benchmarks using the Eleuther AI Language Model Evaluation Harness [19], a widely used framework [20] to test generative language models on a large number of different evaluation tasks.
To further understand the regressive side effects, we conducted a control experiment to compare the model trained to predict tutor responses versus one trained to predict student responses. The results showed that training the LLM on tutor responses did not lead to any performance decline observed when mimicking student responses. This trends highlight that the regressive side effects are a unique challenge specific to training LLMs to replicate student misconceptions.
@@
To counteract the side effects, we propose to incorporate novel start and end hallucination tokens ([hal] and [/hal]) into the LLM training process. These tokens, placed at the beginning and end of each student response, serve as cues to the model, instructing it when to differentiate between providing accurate responses and replicating student misconceptions. Our results indicate a substantial improvement in the model’s performance across all datasets after introducing this token. However, these tokens do not fully restore the model’s baseline performance, underscoring the complexity of the issue.
Through our research, we have brought to the following critical contributions in the realm of personalized education leveraging LLMs:
1. We have uncovered and thoroughly investigated regressive side effects in the LLMs trained for student modeling. This research highlights the paradoxical challenge when LLMs are trained to mimic student misconceptions, potentially compromising their factual integrity and reasoning ability.
2. We have proposed hallucination tokens to mitigate these regressive effects. These tokens, added to the training process, instruct the LLMs to differentiate between mimicking student misconceptions and providing factually accurate responses, substantially improving the model’s performance.
3. Despite the improvements achieved with the hallucination tokens, our research indicates that it does not fully counteract the regressive side effects. This points to the complexity of this issue and underscores the need for further research in this area.
Our research marks a significant step towards understanding the complexities of using LLMs for student modeling. The findings and contributions of this study will fuel further exploration and innovation in this domain, ultimately refining the use of LLMs in personalized learning environments.
2 Related Work
2.1 Student Modeling
Student modeling has long been the cornerstone of personalized learning, with early attempts using rule-based and Bayesian systems to predict student knowledge and behaviors [21]. Recent advancements have shifted towards utilizing machine learning to create more sophisticated models that can adapt to student learning patterns over time [22, 7]. Our work builds upon these foundations by exploring how LLMs can simulate not only the knowledge but also the typical errors and misconceptions students have during the learning process.
2.2 Intelligent Tutoring Systems (ITS)
Intelligent tutoring systems have been designed to provide immediate and personalized instruction or feedback to learners without human intervention [23]. The application of LLMs in ITS presents a novel opportunity to create systems that can engage in more natural and meaningful dialogues with students [3, 12]. Our approach diverges from traditional ITS by focusing on the intentional generation of errors to mimic a student’s learning trajectory, rather than solely providing expert-level instructions [24].
2.3 Large Language Models in Education
The use of LLMs like GPT [5] in education is a relatively new but rapidly growing field of study [25]. These models have been employed for various educational purposes, from generating educational content to serving as conversational agents [26, 12]. However, the challenge of ensuring the truthfulness and reliability of the information provided by LLMs is a recurring concern [27]. Our research contributes to this dialogue by investigating the impact of training LLMs to produce student-like errors and proposing a novel ‘hallucination token’ to manage this trade-off.
2.4 Truthfulness and Reliability in AI
The TruthfulQA benchmark has been instrumental in highlighting the issues of truthfulness in AI-generated content [28]. The ARC challenge further emphasizes the complexity of reasoning required from AI systems beyond simple fact retrieval [29]. Our work is aligned with these challenges, as we seek to understand and improve the truthfulness and reasoning capacity of LLMs when they are trained to replicate student behaviors.
In conclusion, our study intersects with and contributes to the existing body of work in these areas by addressing the unique challenge of training LLMs to authentically mimic student learning processes, including the generation of errors. Our introduction of the “hallucination token” represents a step forward in this domain, suggesting a new direction for future research and development.
@@ 3.1 Data Preparation
The dataset is derived from the CLASS framework [12], which provides a realistic representation of student learning patterns, featuring student misconceptions and the tutor’s rectifications. This dataset provides a rich source of student-tutor dialogues on biology questions sourced from college textbooks.
@@ 4 Experiments and Discussion
Table 1: Performance of Vicuna models on TruthfulQA tasks. The table compares the performance of the original vicuna model, the control model trained to mimic tutor responses in biology (tutor) the model trained to mimic tutor responses in biology (tutor), and the model trained with hallucination tokens (student-hal). The results are presented for three different settings: MC1, MC2, and Generation. MC1 refers to a setting where there is only one correct answer to a question, while MC2 refers to a setting where there are multiple correct answers. For these settings, the performance is measured in terms of accuracy. The generation setting involves the model generating 1-2 sentence answers, with performance evaluated using BLEU and ROUGE scores. The results highlight the significant drop in performance when the model is trained to mimic student responses, demonstrating a regressive side effect in terms of truthfulness. However, the substantial recovery in performance with the introduction of hallucination tokens suggests a promising strategy to mitigate these regressive effects.
@@
In this section, we present our experimental methodology and discuss the findings in detail. The experiments were designed to explore the regressive side effects of training LLMs to mimic student behavior and to assess the effectiveness of our proposed hallucination tokens in mitigating these effects.
@@ 4.1 Experimental Setup
We trained the Vicuna 7B and 13B models [14], one of the best open-source LLMs, on a student-tutor dialogue dataset derived from the CLASS [12] framework. This dataset, which provides a realistic representation of student learning patterns, misconceptions, and the tutor’s rectifications, was used to fine-tune the models to generate outputs that mimic student dialogue. The dataset contains $648$ conversations, which sums up to a total of 20K student-tutor interactions. Average length of conversations is around 400400~{}400400 words, only including the student and tutor fields in the conversation template.
The models were evaluated across seven key benchmarks using the Eleuther AI Language Model Evaluation Harness [19]. These benchmarks include the TruthfulQA [16], ARC [15], HellaSwag [30], Winogrande [31], MMLU [32], HaluEval Dialogue [17], and MemoTrap [18]. Each of these benchmarks tests different aspects of the model’s performance, including its truthfulness, reasoning abilities, ability to recognize hallucinations, and memory-based task performance.
@@ 4.2 In-depth Analysis: TruthfulQA
TruthfulQA is a benchmark specifically designed to measure the truthfulness of a language model’s responses across a wide range of categories. It tests the model’s ability to avoid generating false answers learned from imitating human texts, a challenge that is particularly relevant to our study. Given the importance of truthfulness in educational contexts and the unique challenges posed by training models to mimic student misconceptions, we believe that a rigorous analysis of our models’ performance on TruthfulQA is warranted.
In this section, we present our findings from the TruthfulQA benchmark, exploring the impact of training models to mimic student behavior and the effectiveness of our proposed hallucination tokens in mitigating any negative effects. We delve into the results from the multiple-choice and generation tasks within TruthfulQA, providing a comprehensive view of our models’ truthfulness in different contexts.
@@
TruthfulQA Multiple-Choice Setting 2 (MC2) Findings. In the second multiple-choice setting, where multiple correct labels are possible, we observed a similar trend to the MC1 setting. The student-7b model experienced a significant drop in accuracy, from 50.37% in the vicuna-7b model to 36.14% when trained to mimic student responses. However, the introduction of hallucination tokens led to a notable improvement in performance, with the student-7b model’s accuracy recovering to 44.68%.
@@
Table 2: Comparative performance of Large Language Models (LLMs) on various benchmarks before and after the introduction of hallucination tokens, with a control experiment involving tutor models. The table presents the performance of Vicuna 7B models across five key benchmarks: ARC Reasoning, Hallucination Evaluation Dialogue (HaluDial), Hallucination Memorization Trap (MemoTrap), TruthfulQA (TQA), HellaSwag (HSwag), MMLU, and Winogrande (WinoG). The numbers in parentheses (e.g., 25-S in ARC) represent the number of few-shot examples provided to the model during evaluation. The performance is measured in terms of accuracy percentage. The table compares the performance of the original vicuna models, tutor models, student models, and student models trained with hallucination tokens (student-hal). The results highlight the significant drop in performance when the model is trained to mimic student responses, demonstrating regressive side effects across multiple tasks. However, the introduction of hallucination tokens leads to a substantial recovery in performance across all benchmarks, underscoring their potential in mitigating these regressive effects.

### v2

# Student Data Paradox and Curious Case of Single Student-Tutor Model: Regressive Side Effects of Training LLMs for Personalized Learning

@@ Abstract
The pursuit of personalized education has led to the integration of Large Language Models (LLMs) in developing intelligent tutoring systems. To better understand and adapt to individual student needs, including their misconceptions, LLMs need to be trained on extensive datasets of student-tutor dialogues. Our research uncovers a fundamental challenge in this approach: the “Student Data Paradox.” This paradox emerges when LLMs, trained on student data to understand learner behavior, inadvertently compromise their own factual knowledge and reasoning abilities. We investigate this paradox by training state-of-the-art language models on student-tutor dialogue datasets and evaluating their performance across multiple benchmarks. These benchmarks assess various aspects of language model capabilities, including reasoning, truthfulness, and common sense understanding. Our findings reveal significant declines in the models’ performance across these diverse benchmarks, indicating a broad impact on their capabilities when trained to model student behavior. Our research makes two primary contributions: (1) empirical demonstration of the Student Data Paradox through quantitative analysis of model performance, and (2) introduction of “hallucination tokens” as a mitigation strategy. These tokens, while improving performance, highlight the persistent challenge of balancing accurate student behavior modeling with maintaining the LLM’s integrity as an educational tool. This study emphasizes the need for innovative solutions to reconcile the conflicting goals of faithfully understanding diverse student cognition while preserving the model’s ability to provide accurate information and guidance.
Student Data Paradox and Curious Case of Single Student-Tutor Model:
Regressive Side Effects of Training LLMs for Personalized Learning
Shashank Sonkar, Naiming Liu, Richard G. Baraniuk Rice University shashank.sonkar@rice.edu
@@ 1 Introduction
Personalized education is undergoing a significant transformation, driven by the rapid advancement of Large Language Models (LLMs) Jurenka et al. (2024). These AI systems are being developed to serve as adaptive tutors, capable of understanding and responding to individual student needs OpenAI (2024); Liu et al. (2023); Sonkar et al. (2024c). The incorporation of student data into the training process is an important step towards creating truly personalized learning experiences. By training LLMs on diverse student data, including dialogues, errors, and problem-solving approaches, these models can learn to identify knowledge gaps, recognize common misconceptions, and provide personalized support tailored to each learner’s unique needs. However, this approach is at odds with traditional LLM training, where data quality is of utmost importance Goldberg (2022); Samaniego et al. (2022). Student data, by nature, is prone to errors and misconceptions, which raises concerns about the impact on the LLMs’ factual knowledge and reasoning abilities. This tension between the need to understand student misconceptions accurately and the imperative to maintain the model’s factual integrity creates a paradoxical challenge, which we term the “Student Data Paradox.
In this paper, we study this timely topic: what happens when we train an LLM on student-tutor interaction data? Our research reveals that while training LLMs on student data can enable models to accurately simulate authentic student behaviors and misconception, it comes at a significant cost – LLMs’ own factual integrity and reasoning capabilities are compromised. This paradox poses a serious concern, as the primary purpose of any educational model is to provide accurate and reliable information to learners.
Figure 1: Training LLMs on student-tutor dialogues to model student behavior: an example from the CLASS dataset. This approach creates “Student Data Paradox”: while the model accurately simulates student responses and misconceptions, its own factual knowledge and reasoning abilities are compromised. We introduce hallucination tokens ([hal], [/hal]) during training to instruct the model when to replicate misconceptions versus provide accurate information, showing promise in mitigating the paradox.
We investigated this paradox by fine-tuning Llama Touvron et al. (2023) and Vicuna models Chiang et al. (2023), on the CLASS dataset Sonkar et al. (2023) - a collection of student-tutor dialogues on college-level biology questions (example in figure 1). Training LLMs to simulate student behavior significantly decreased the models’ performance across various benchmark datasets, including the ARC reasoning challenge, TruthfulQA, Hallucination Evaluation Dialogue, and MemoTrap Clark et al. (2018a); Lin et al. (2022); Li et al. (2023); McKenzie et al. (2022). We present a detailed analysis across nine key benchmarks using the Eleuther LLM leaderboard Gao et al. (2023) to provide a thorough assessment of LLMs’ reasoning and general knowledge capabilities. In contrast, a control experiment revealed that training LLMs to simulate tutor behavior did not lead to similar performance declines. This finding highlights that the observed regressive effects are uniquely associated with training LLMs to replicate student misconceptions.
@@
A seemingly intuitive solution to the paradox would be to create separate student and tutor models. However, we present a formal theorem demonstrating that this approach necessitates potentially vast number of student models, making it impractical. This finding emphasizes the need for a unified framework to capture both accurate knowledge and diverse student misconceptions. To this end to counteract the side effects, we propose to incorporate novel start and end hallucination tokens ([hal] and [/hal]) into the LLM training process. These tokens, placed at the beginning and end of each student response, serve as cues to the model, instructing it when to differentiate between providing accurate responses and replicating student misconceptions. Our results indicate a substantial improvement in the model’s performance across all datasets after introducing this token. However, these tokens do not fully restore the model’s baseline performance, underscoring the complexity of the issue.
Contributions: Our research makes important contributions to the use of LLMs in modeling student behavior for adaptive tutoring systems. We identify and study the Student Data Paradox in LLMs trained on student-tutor interactions, revealing a trade-off between accurately modeling student misconceptions and maintaining the LLM’s factual integrity and reasoning ability (section 3 and  4). Through a formal theorem discussed in section 2, we show that the intuitive solution of separating student and tutor models is fundamentally flawed, necessitating a more sophisticated, unified approach. We propose hallucination tokens as a method for this unified approach. These tokens, integrated into the training process, enable LLMs to differentiate between modeling student misconceptions and providing accurate information, significantly improving model performance.
Our study exposes the critical trade-offs in training LLMs on student data and offers a concrete solution with hallucination tokens. This work lays the foundation for developing more robust and effective AI tutors that can accurately model student behavior without compromising factual integrity.
2 Theorem: The Case for a Single Student-Tutor Model
To address the “Student Data Paradox”, one may propose a seemingly straightforward solution - the creation of two distinct models: a student model and a tutor model. This approach seems intuitive: one model to capture diverse student behaviors and misconceptions, and another to represent the ideal tutor with perfect knowledge. However, this seemingly straightforward solution harbors significant complexities and challenges that are not immediately apparent. Our goal is to show that what may appear to be a simple dichotomy between student and tutor models necessarily evolves into a multi-model approach with potentially multitude of student models.
To illustrate the complexity of this problem, consider a simple algebraic equation: $Ax=B$. A tutor (or a student with perfect knowledge) would correctly solve this as $x=B/A$. However, different students might have various misconceptions, leading to incorrect solutions such as $x=A/B$, $x=B-A$, or $x=B+A$. The proposal to separate tutor and student models aimed to resolve the inconsistency between correct and incorrect knowledge. However, this solution overlooks an important issue: the inconsistencies among the incorrect rules themselves. These mutually contradictory misconceptions cannot be accurately represented by a single student model without compromising the model’s internal consistency. Consequently, to faithfully represent the diverse range of student misconceptions, we would need multiple student models, each consistently representing a different set of misconceptions. This realization demonstrates that the apparent dichotomy between a tutor model and a student model inevitably evolves into a multi-model scenario, with potentially numerous student models. Thus, the initial attempt to resolve inconsistencies by separating correct and incorrect knowledge actually leads to a proliferation of models, highlighting the need for a more sophisticated, unified approach to effectively capture both accurate knowledge and diverse student misconceptions.
The complexity illustrated in the previous example underscores the need for a more rigorous, formal treatment of the problem. To proceed with our analysis, we must first establish clear definitions and state our working hypothesis. These will form the foundation for our main theorem and subsequent proofs, allowing us to formally demonstrate why the apparent dichotomy between student and tutor models inevitably leads to a multi-model scenario.
2.1 Definitions
Definition 1.
Let $R$ be the set of all possible rules or misconceptions a student might have.
Definition 2.
Two rules $r_{1},r_{2}\in R$ are considered logically consistent if and only if there exists a possible state of the world in which both $r_{1}$ and $r_{2}$ can be simultaneously true or applicable.
Definition 3.
For any $R_{i},R_{j}\subseteq R$, let $C(R_{i},R_{j})$ be a function that returns true if all rules in $R_{i}$ are logically consistent with all rules in $R_{j}$, and false otherwise.
Definition 4.
Let $S$ be the set of all distinct student models, where each model $M_{i}\in S$ represents a unique, internally consistent set of rules or misconceptions. Formally, $S=\{M_{i}:M_{i}\text{ represents a maximal set }R_{i}\subseteq R\text{ such that }\forall r_{1},r_{2}\in R_{i},C(\{r_{1}\},\{r_{2}\})=true\}$ where $R$ is the set of all possible rules or misconceptions.
Definition 5.
A "perfect student" is defined as one who consistently provides correct answers and possesses comprehensive subject understanding, functionally equivalent to a tutor in terms of knowledge representation.
Hypothesis 1.
Models struggle to simultaneously represent multiple logically inconsistent rules while maintaining internal consistency.
If this hypothesis is false, it implies that a single student-teacher model is theoretically feasible and supports our unified approach of a single student-tutor model.
2.2 Main Theorem
Theorem 1.
The apparent dichotomy between a student model and a tutor model necessarily evolves into a multi-model approach with potentially numerous student models.
Proof.
1. Assume, for contradiction, that a single student model $M_{s}$ can accurately represent all student misconceptions.
2. Consider two sets of misconceptions $R_{1},R_{2}\subset R$ such that $C(R_{1},R_{2})=false$.
3. By our hypothesis, $M_{s}$ cannot consistently represent both $R_{1}$ and $R_{2}$ simultaneously.
4. Therefore, we need at least two distinct student models, contradicting our assumption of a single student model.
5. Let $n$ be the number of maximal subsets of $R$ containing mutually consistent rules.
6. We require $n$ distinct models to represent all sets of consistent misconceptions.
7. Let $R_{c}\subset R$ be the set of correct rules. One of these models, $M_{c}$, corresponds to $R_{c}$, representing both a perfect student and a tutor.
8. Therefore, the total number of necessary models is $n$, where $n\geq 2$ and one model serves as both the perfect student and tutor model. ∎
2.3 Corollaries
Corollary 1.
The number of required student models is at least equal to the number of maximal sets of mutually consistent misconceptions.
Corollary 2.
One model in the set of student models is functionally equivalent to the tutor model.
Corollary 3.
The apparent two-model approach (student and tutor) actually requires $n$ models, where $n$ can be large, with one of these models serving as both the perfect student and tutor model.
This formal treatment demonstrates that the intuitive separation of student and tutor models is insufficient to address the complexities of representing diverse student misconceptions. The proliferation of necessary models to accurately capture different sets of consistent misconceptions reveals the impracticality of this approach. These findings point to the necessity of a unified approach to model training that can efficiently represent both accurate knowledge and diverse student misconceptions within a single framework. Such an approach would need to address the challenge of maintaining internal consistency while capturing a wide range of potentially conflicting rules.
In the following section, we propose a novel solution to address this challenge: the incorporation of hallucination tokens into the training process of LLMs. This approach aims to provide a unified framework that can represent both accurate knowledge and diverse student misconceptions without resorting to multiple, separate models.
@@ 3.1 Data Preparation
The dataset is derived from the CLASS framework Sonkar et al. (2023), which provides a realistic representation of student learning patterns, featuring student misconceptions and the tutor’s rectifications. This dataset provides a rich source of student-tutor dialogues on biology questions sourced from college textbooks.
@@ 4 Experiments and Discussion
Table 1: Performance of Vicuna models on TruthfulQA tasks. The table compares the performance of the original vicuna model, the control model trained to model tutor responses in biology (tutor) the model trained to model tutor responses in biology (tutor), and the model trained with hallucination tokens (student-hal). The results are presented for three different settings: MC1, MC2, and Generation. MC1 refers to a setting where there is only one correct answer to a question, while MC2 refers to a setting where there are multiple correct answers. For these settings, the performance is measured in terms of accuracy. The generation setting involves the model generating 1-2 sentence answers, with performance evaluated using BLEU and ROUGE scores. The results highlight the significant drop in performance when the model is trained to model student responses, demonstrating a regressive side effect in terms of truthfulness. However, the substantial recovery in performance with the introduction of hallucination tokens suggests a promising strategy to mitigate these regressive effects.
@@
In this section, we present our experimental methodology and discuss the findings in detail. The experiments were designed to explore the regressive side effects of training LLMs to model student behavior and to assess the effectiveness of our proposed hallucination tokens in mitigating these effects.
@@ 4.1 Experimental Setup
We trained the Vicuna 7B and 13B models Chiang et al. (2023), one of the best open-source LLMs, on a student-tutor dialogue dataset derived from the CLASS Sonkar et al. (2023) framework. This dataset, which provides a realistic representation of student learning patterns, misconceptions, and the tutor’s rectifications, was used to fine-tune the models to generate outputs that model student dialogue. The dataset contains $648$ conversations, which sums up to a total of 20K student-tutor interactions. Average length of conversations is around 400400~{}400400 words, only including the student and tutor fields in the conversation template.
The models were evaluated across seven key benchmarks using the Eleuther AI Language Model Evaluation Harness Gao et al. (2023). These benchmarks include the TruthfulQA Lin et al. (2022), ARC Clark et al. (2018a), HellaSwag Zellers et al. (2019), Winogrande Sakaguchi et al. (2019), MMLU Hendrycks et al. (2020), HaluEval Dialogue Li et al. (2023), and MemoTrap McKenzie et al. (2022). Each of these benchmarks tests different aspects of the model’s performance, including its truthfulness, reasoning abilities, ability to recognize hallucinations, and memory-based task performance.
@@ 4.2 In-depth Analysis: TruthfulQA
TruthfulQA is a benchmark specifically designed to measure the truthfulness of a language model’s responses across a wide range of categories. It tests the model’s ability to avoid generating false answers learned from imitating human texts, a challenge that is particularly relevant to our study. Given the importance of truthfulness in educational contexts and the unique challenges posed by training models to model student misconceptions, we believe that a rigorous analysis of our models’ performance on TruthfulQA is warranted.
In this section, we present our findings from the TruthfulQA benchmark, exploring the impact of training models to model student behavior and the effectiveness of our proposed hallucination tokens in mitigating any negative effects. We delve into the results from the multiple-choice and generation tasks within TruthfulQA, providing a comprehensive view of our models’ truthfulness in different contexts.
@@
TruthfulQA Multiple-Choice Setting 2 (MC2) Findings. In the second multiple-choice setting, where multiple correct labels are possible, we observed a similar trend to the MC1 setting. The student-7b model experienced a significant drop in accuracy, from 50.37% in the vicuna-7b model to 36.14% when trained to model student responses. However, the introduction of hallucination tokens led to a notable improvement in performance, with the student-7b model’s accuracy recovering to 44.68%.
@@
Table 2: Comparative performance of Large Language Models (LLMs) on various benchmarks before and after the introduction of hallucination tokens, with a control experiment involving tutor models. The table presents the performance of Vicuna 7B models across five key benchmarks: ARC Reasoning, Hallucination Evaluation Dialogue (HaluDial), Hallucination Memorization Trap (MemoTrap), TruthfulQA (TQA), HellaSwag (HSwag), MMLU, and Winogrande (WinoG). The numbers in parentheses (e.g., 25-S in ARC) represent the number of few-shot examples provided to the model during evaluation. The performance is measured in terms of accuracy percentage. The table compares the performance of the original vicuna models, tutor models, student models, and student models trained with hallucination tokens (student-hal). The results highlight the significant drop in performance when the model is trained to model student responses, demonstrating regressive side effects across multiple tasks. However, the introduction of hallucination tokens leads to a substantial recovery in performance across all benchmarks, underscoring their potential in mitigating these regressive effects.